Time to Swing:
The period T (time in seconds for one complete cycle) of a simple pendulum is related to the L (in feet) of teh pendulum by the formula 8T^2=pie^2L. If a child is on a swing with a 10-foot chain, then how long does it take to complete one cycle of the swing.
This problem is really throwing me off and it has a lot to do with the "pie." Any suggestions?

Answer by stanbon(75887)

The period T (time in seconds for one complete cycle) of a simple pendulum is related to the L (in feet) of teh pendulum by the formula 8T^2=pie^2L. If a child is on a swing with a 10-foot chain, then how long does it take to complete one cycle of the swing.
Your formula is confusing.
It could mean (pi)(e^2)L or (pi)^(2L) or (pi^2)L
Assuming it means the following you should get:
8T^2=(pi^2)L
8T^2=(3.1416...^2)(10)
T^2 = (1/8)(9.86...)(10)
T^2=(1/8)(98.6...)
T^2= 12.337...
T= 3.51 seconds
Cheers,
Stan H.